Two light spotted the same mark on the tower, the mark was 36 ft. from the ground. The first light forming 97 ft. ray and 20º angle from the ground and the second one nearer to the mark was 70 ft. from the first. Find the length of the ray from the second light to the mark


1
Expert's answer
2021-10-25T16:53:02-0400

"AB=\sqrt{97^2-36^2}=90.072ft"

"x=90.072ft-70ft=20.072ft"

length of second ray to the mark"=\sqrt{36^2+20.072^2}=41.22ft"